Suppose the following demand and supply function:

Qd = 750 – 25P

Qs = -300 + 20 P


       i.           Find equilibrium price and quantity

     ii.           Find consumer and producer surplus



1
Expert's answer
2021-06-01T13:33:21-0400

Answer:-

Qd = 750 – 25P

Qs = -300 + 20 P

i.

Equilibrium arises where:

Qs= Qd

-300+20P= 750-25P

20P+25P= 750+300

45P= 1050

P*= "1050\over45" = 23.33 Equilibrium price

Q*= -300+20(P*)= -300+20(23.33)= -300+466.6= 166.6 Equilibrium quantity

ii.

Qd = 750 – 25P

When Qd=0, Pd= 30

Consumer surplus= ("1\over2" )(Pd-P*)(Q*)= (1/2)(30-23.33)(166.6)= 555.611

Qs = -300 + 20 P

When Qs=0, Ps= 15

Producer Surplus= ("1\over2" )(P*-Ps)(Q*)= (1/2)(23.33-15)(166.6)= 693.889